Mole fraction of a solute in benzene solvent is 0.2. The molality of the solution will be??

Grade:12

1 Answers

Arun
6 years ago
Dear Ashutosh
mole fraction = 0.2.therefore in 100 moles, 20 moles are of solute and 80 moles are of benzene(solvent)mass of benzene = 80 * Mol.Wt of Benzene = 80* 78 = 6240gfor 6240 g of solvent we have 20 moles of soluteTherefore for 1000g of solvent we have (20)*(1000)/(6240) = 3.20 molal RegardsArun (askIITians forum expert)